### v2

# FedCQA: Answering Complex Queries on Multi-Source Knowledge Graphs via Federated Learning

Abstract.
Complex logical query answering is a challenging task in knowledge graphs (KGs) that has been widely studied. The ability to perform complex logical reasoning is essential and supports various graph reasoning-based downstream tasks, such as search engines. Recent approaches are proposed to represent KG entities and logical queries into embedding vectors and find answers to logical queries from the KGs. However, existing proposed methods mainly focus on querying a single KG and cannot be applied to multiple graphs. In addition, directly sharing KGs with sensitive information may incur privacy risks, making it impractical to share and construct an aggregated KG for reasoning to retrieve query answers. Thus, it remains unknown how to answer queries on multi-source KGs. An entity can be involved in various knowledge graphs and reasoning on multiple KGs and answering complex queries on multi-source KGs is important in discovering knowledge cross graphs. Fortunately, federated learning is utilized in knowledge graphs to collaboratively learn representations with privacy preserved. Federated knowledge graph embeddings enrich the relations in knowledge graphs to improve the representation quality. However, these methods only focus on one-hop relations and cannot perform complex reasoning tasks. In this paper, we apply federated learning to complex query-answering tasks to reason over multi-source knowledge graphs while preserving privacy. We propose a Federated Complex Query Answering framework (FedCQA), to reason over multi-source KGs avoiding sensitive raw data transmission to protect privacy. We conduct extensive experiments on three real-world datasets and evaluate retrieval performance on various types of complex queries. The experimental results show that our proposed methods can effectively retrieve answers to cross knowledge graph queries while keeping raw data secret in local devices.
privacy preserving, knowledge graph, complex query answering, federated Learning
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacy Privacy protectionsccs: Information systems Data mining
1. Introduction
Knowledge graphs (KG) are structured knowledge databases representing external knowledge using triplets and describing the relationships between entities. KG plays an important role and has been widely applied in numerous real-world applications, such as recommender systems (Wang et al., 2019; Cao et al., 2019) and search engines (Xiong et al., 2017). Among them, complex query answering (CQA) is an important but challenging task in knowledge graph reasoning and can be used for supporting various downstream tasks (Bai et al., 2023c; Ren et al., 2023). CQA aims to retrieve answers that satisfy given logical expressions (Hamilton et al., 2018; Ren et al., 2020), which are often defined in predicate logic forms with relation projection operations, existential quantifiers $\exists$, logical conjunctions $\wedge$, disjunctions $\vee$, etc. For example in Figure 1, given a logical query $q$, We aim to find all the research topic entities $V_{?}$ for which there exists Nobel Prize winners $V$ who were born in Germany and conducted studies in that specific field. CQA is challenging due to the incompleteness of knowledge graphs and the complexity of query matching (Bai et al., 2023c). To solve these problems, query encoding methods are commonly used in CQA. These methods try to parse complex queries into computation graphs and parameterize both queries and entities using different neural networks simultaneously. After the encoding process, these methods retrieve answers according to the similarity scores between the parameterized embeddings of answer sets and queries  (Hamilton et al., 2018; Ren et al., 2020).
Existing query encoding methods have gained remarkable success in addressing complex query answering tasks. However, these query answering models are limited to the utilization of a single central knowledge graph and cannot be extended to multiple KGs. With the development of knowledge graphs, KGs have grown significantly in size and encompass knowledge from various domains. Consequently, constructing a knowledge graph that includes all related entities and relations has become difficult and it is impractical to access a central KG with all the triples needed (Peng et al., 2021a; Chen et al., 2021; Zhang et al., 2022). Collaborations between various knowledge graph holders are helpful to answer more complicated queries. For example, as shown in Figure 1, there are multiple knowledge graphs from different domains. A complex query may consist of entities and relations from multiple knowledge graphs, preventing a local query answering model on a single knowledge graph from answering that cross-graph queries. However, there are various reasons hindering the data sharing between knowledge graph holders. For example, the growing attention on privacy, regulations such as the General Data Protectioan Regulation (GDPR), and commercial interest between data holders, etc.
Figure 1. An example of cross knowledge graph queries on distributed knowledge graphs. The relations and entities in a KG complex query can be from multiple knowledge graphs which cannot be solved in a single local knowledge graph.
To solve the above challenges, federated learning (McMahan et al., 2017; Yang et al., 2019) has been proposed which allows multiple participants to distributedly train a global model collaboratively without raw data transmission. Federated learning has been widely applied in various domains (Long et al., 2020), including knowledge graph embedding. In such systems, raw knowledge graph triplets are kept on the local devices, participants train local models, and only gradient will be transferred to learn a global knowledge graph embedding model (Chen et al., 2021; Zhang et al., 2022; Peng et al., 2021a). In the training process, homomorphic encryption (HE) (Paillier, 1999), secure multi-party computation (SMPC) (Mohassel and Zhang, 2017), and differential privacy (Dwork, 2008; Geyer et al., 2017) are widely applied in federated learning to improve security. FedAvg (McMahan et al., 2017) is a commonly used technique in federated learning that updates global models by averaging local models trained on individual data and has been widely studied in various areas, including federated knowledge graph embedding (Chen et al., 2021; Zhang et al., 2022). Although federated learning can protect raw data, recent studies have indicated that learned representations can still potentially leak privacy (Hu and Song, 2023; Duddu et al., 2020), where an attacker can infer sensitive information from the learned embeddings.
While federated learning has been widely applied in knowledge graph embeddings, existing federated knowledge graph embedding works only focus on learning high quality representations for simple downstream tasks, such as knowledge graph completion (Chen et al., 2021; Zhang et al., 2022; Huang et al., 2022; Tang et al., 2023), and lacks the ability of reasoning over knowledge graphs and retrieving answers to complex queries. To solve the constraints, we propose a Federated Complex Query Answering framework (FedCQA), to reason over multi-source KGs avoiding sensitive raw data transmission to protect data privacy. FedCQA can be applied to different central complex query answering models. It leverages federated learning techniques to train local query answering models in local knowledge graphs and align knowledge graph embeddings for global queries. Different from other federated knowledge graph embedding models, FedCQA server not only takes the responsibility of aggregating global models, but decomposing given complex queries to sub-queries to compute the query encoding and retrieve answers from distributed knowledge graphs under the protection of multi-party computation to avoid global model storage. Meanwhile, to better evaluate federated complex query answering systems’ performance, we create a benchmark on three widely used knowledge graph datasets. We evaluate our proposed FedCQA on the benchmark and the experiment results show the effectiveness of retrieving answers to complex queries from multi-source knowledge graphs. We summarize our major contributions as follows:
  • To the best of our knowledge, we are the first to extend federated knowledge graph embedding systems to complex query answering tasks, which is critical for knowledge graph holders’ collaboration.
  • Based on three public datasets, we propose a benchmark for evaluating the retrieval performance of federated complex query answering systems. The benchmark systematically evaluates the retrieval performance facing cross knowledge graph queries.
  • We propose FedCQA, a federated complex query answering system that can retrieve answers to complex queries from distributed knowledge graphs with privacy preserved. Extensive experiments conducted on three datasets demonstrate its high performance facing cross knowledge graphs queries.
The rest of the paper is organized as follows. We review the related works in Section 2. Section 3 introduces the preliminary and definition of the distributed complex query answering systems. Section 4 introduces the framework of FedCQA in detail. Section 5 evaluates the performance of FedCQA on the benchmark based on three real-world datasets. Finally, we make a conclusion of our work in Section 6.
2. Related Work
In this section, we summarize the related work. Our work is closely related to Complex query Answering (CQA) and federated knowledge graph embedding.
2.1. Complex Query Answering
Complex query answering (CQA) is a crucial task in knowledge graphs, where a model is learned to process and answer complex logical queries based on knowledge graph reasoning. Query encoding is effective and efficient for complex query answering. These methods represent complex queries into various structures and effectively search answers among candidate knowledge graph entities. GQE (Hamilton et al., 2018), Q2B (Ren et al., 2020) and HypeE (Liu et al., 2021) encode queries to vectors, hyper-rectangle and hyperbolic embeddings, respectively. To support negation operators, various encoding methods are proposed: Q2P (Bai et al., 2022) and ConE (Zhang et al., 2021) use multiple vectors to represent queries. BetaE (Ren and Leskovec, 2020), GammaE (Yang et al., 2022), PERM (Choudhary et al., 2021) propose to use various probabilistic distributions to encode complex logic graph queries. Some methods, like LMPNN (Wang et al., 2022) CQD (Arakelyan et al., 2020), Var2Vec (Wang et al., 2023a) take pre-train embeddings on simple link prediction tasks and apply logic operators to answer complex queries. Meanwhile, Neural structures are utilized to encode complex queries: BiQE (Kotnis et al., 2021) and KgTransformer (Liu et al., 2022) are proposed to use transformer, SQE (Bai et al., 2023c) applies sequential encoders, GNN-QE (Zhu et al., 2022) and StarQE (Alivanistos et al., 2021) use message-passing graph neural networks to encode queries, respectively. There are also encoding methods proposed to encode various knowledge graph types: NRN (Bai et al., 2023b) is proposed to encode numerical values, MEQE (Bai et al., 2023a) extends logical queries over events, states, and activities.
While there are numerous existing complex query answering methods, these methods mainly focus on a single large knowledge graph. There are some methods proposed to reason over multi-view and temporal and varying knowledge graphs: MORA (Xi et al., 2022) ensembles multi-view knowledge graphs to scale up complex query answering, TTransE (Leblay and Chekol, 2018) and TRESCAL (Wang et al., 2015) can be applied on temporal knowledge graphs. However, these methods need raw data transmission and privacy protection is not considered. Conducting privacy-preserving complex query answering on multi-source knowledge graphs is still unexplored. With the growing attention on privacy and data protection, sensitive data is unable to freely circulate among data holders, and complex query answering is forced to be conducted on multiple knowledge graphs collaboratively. Our research introduces federated learning to existing complex query answering so that we can apply reasoning over distributed knowledge graphs without raw data sharing.
2.2. Federated Knowledge Graph Embedding
In recent years, federated learning has emerged as a promising approach to address privacy and scalability concerns in machine learning. It allows data owners to participate in model co-construction without raw data transmission to reduce privacy leakage risks (McMahan et al., 2017; Yang et al., 2019). Various studies have been conducted to explore the potential of federated learning in different domains, such as recommender system (Yang et al., 2020), finance (Long et al., 2020) etc. Federated knowledge graph embedding is one of the domains in which federated learning has been widely applied. FedE (Chen et al., 2021) learns knowledge graph embeddings locally and aggregates all local models in a global server for higher representation quality. FedR (Zhang et al., 2022) proposes to learn representation with privacy-preserving relation aggregation to avoid privacy leakage risks in entity embedding and reduce communication costs. FedCKE (Huang et al., 2022), FedMKGC (Tang et al., 2023) extend federated learning to learn global representations from different domains and multilingual knowledge graphs. FedEC (Chen et al., 2022b) applies contrastive learning to tackle data heterogeneity in knowledge graphs. MaKEr (Chen et al., 2022a) and MorseE (Chen et al., 2022c) utilize meta-learning to transfer knowledge among knowledge graphs to train graph neural networks for unseen knowledge extrapolation and inductive learning. DP-FLames (Hu et al., 2023) quantifies the privacy threats and incorporates private selection in federated knowledge embeddings. FLEST (Wang et al., 2023b) decomposes the embedding matrix and enables the sharing of latent representations to reduce the risks of privacy leakage and communication costs, FedM (Hu et al., 2022b) splits the duty of aggregating entities and relations to reduce the risks of graph reconstruction attacks. FKGE (Peng et al., 2021b) applies differential privacy and avoids the need for a central server.
While existing federated knowledge graph embedding methods are proposed to distributedly learn high quality representations with privacy preservation, there are some works indicate that the learned embeddings are informative and vulnerable to various privacy attacks (Hu et al., 2022a), even in federated scenarios (Hu and Song, 2023; Hu et al., 2023). Besides, these methods all lack the ability to answer complex queries on multi-source knowledge graphs which is critical for more complicated downstream tasks. Our research expands the simple federated knowledge graph embedding to answer complex queries on distributed knowledge graphs.
3. Preliminary and Problem Formulation
3.1. Preliminary
[…]
3.2. Complex Logical Query
The complex knowledge graph query is defined in existential positive first-order logic form, consisting of various types of logic expressions like existential quantifiers $\exists$, logic conjunctions $\wedge$, and disjunctions $\vee$. In the logical expression, there is a set of anchor entities $V_{a}\in\mathcal{V}$ denotes given context, existential quantified variables $V_{1},V_{2},...,V_{k}\in\mathcal{V}$, and a unique variable $V_{?}$ denotes our query target. The complex query intends to find the target answers $V_{?}\in\mathcal{V}$, such that there are $V_{1},\cdots,V_{k}\in\mathcal{V}$ in the knowledge graph that can satisfy the given logical expressions simultaneously. Following the definition in (Ren et al., 2020), the complex query expression can be converted to the disjunctive normal form (DNF) in the following:
(1) $\begin{split}q[V_{?}]&=V_{?}.V_{1},...,V_{k}:c_{1}\lor c_{2}\lor...\lor c_{n}\\ c_{i}&=e_{i,1}\land e_{i,2}\land...\land e_{i,m},\end{split}$
where $e_{i,j}$ is the atomic logic expression, which can be the triplet $(V,r,V^{\prime})$ denotes relation $r$ between entities $V$ and $V^{\prime}$, $c_{i}$ is the conjunction of several atomic logic expressions $e_{i,j}$. $V,V^{\prime}$ are either anchor entities or existentially quantified variables.
Figure 2. The training and retrieval process of FedCQA. The blue line denotes the training process, and the green line denotes the retrieval process. In the training, clients’ CQA models are trained on respective knowledge graphs. At each round, local CQAs are aggregated at the server and updated using the global parameters. Among them, the embeddings are protected by secret aggregation so that the server can not access them. In the retrieval, each query is decomposed into sub-queries. Clients compute sub-query embeddings which the server are used to aggregate query embeddings. Answer scores are computed at clients and are aggregated at server to retrieve answers.
3.3. Knowledge Graph Set Query
Knowledge graphs are owned by different data holders and cannot be shared directly with each other, therefore, a complex query $q$ can involve entities and relations from different knowledge graphs. We define those queries as follows:
Definition 3.0 (Cross-graph Query).
A complex query $q$ is a cross-graph query if there exists query answers $V_{?}\in\mathcal{V}$, such that there are $V_{1},\cdots,V_{k}\in\mathcal{V}$ in the knowledge graph that can satisfy the given logical expressions and the atomic expressions in the query can not be found in a single knowledge graph.
For example in figure 1, the entity ”Physics” is the answer to the query $q$ because there exists an existentially quantified variable ”Einstein” that can satisfy the logical expression. The query $q$ is a cross-graph query as the atomic expressions in the query are from different knowledge graphs and can not be found in a single graph. For example, Win(Einstein, Novel Prize) and BornIn(Einstein, Germany) are two atomic expressions from different knowledge graphs. When the distributed knowledge graph set faces cross-graph queries, the answer can not be inferred from one single knowledge graph. Besides, We have the definition of in-graph query correspondingly:
Definition 3.0 (In-graph Query).
A complex query $q$ is an in-graph query if for all answers $V_{?}\in\mathcal{V}$ to the query, such that there are $V_{1},\cdots,V_{k}\in\mathcal{V}$ in the knowledge graph that can satisfy the given logical expressions and the atomic expressions in the query are from a single knowledge graph.
These in-graph queries can retrieve answers according to one of the knowledge graphs in the KG set and can be solved with existing complex query answering models. However, in a distributed knowledge graph complex query answering system, the in-graph queries may retrieve more answers as more knowledge are provided.
3.4. Problem Formulation
Given a knowledge graph set $\mathcal{G}=\{g_{1},...,g_{N}\}$ with $N$ knowledge graphs. Every knowledge graph is owned by independent data holders and can not be shared to construct a unified knowledge graph, which means that the triples in each knowledge graph will stay private at local devices. The knowledge graphs in the $\mathcal{G}$ are related and have part of entities and relation overlapped. There are complex queries involving elements from the knowledge graph set $\mathcal{G}$ and can be classified as cross-graph queries and in-graph queries. We aim to construct a distributed complex query answering system to reason over multi-source knowledge graphs and retrieve answers to the queries while keeping privacy preserved. To achieve this, we assume that there is an honest but curious server managing the federated complex query answering system. Because the learned embeddings are vulnerable to various privacy attacks, the embeddings on the server should be further protected.
4. Federated Complex Query Answering
In this section, we introduce the learning and query retrieval process of our proposed FedCQA.
4.1. Model Learning
We first introduce the training of FedCQA. As shown in Figure 2, FedCQA has a central server and a set of clients. Each client has a knowledge graph with overlapping entities of others. The server takes the responsibility of aggregating parameters and organizing the training and retrieval process. The clients train a local CQA model based on their knowledge graph. According to the sensitivity of the parameters, we divide the query encoding methods into two parts: operator function with parameter $\Theta$ and entity embeddings $\mathbf{E}$. For the operator function, the client directly sends the parameter to the server for aggregation and receives the global function to update the local operator function, which is similar to FedAvg (McMahan et al., 2017). Therefore, in the following parts, we only introduce the entity embeddings aggregation in FedCQA.
4.1.1. Secret Aggregation
There are various techniques, like homomorphic encryption (HE) (Paillier, 1999), secure multi-party computation (MPC) (Mohassel and Zhang, 2017), and differential privacy (Dwork, 2008; Geyer et al., 2017) to protect the uploaded parameters, however the protection of aggregated global model are often ignored. Unfortunately, the global model is informative and vulnerable to privacy attacks  (Zhang et al., 2022). Hence we propose a secret aggregation applied in parameter aggregation which can prevent the global server from knowing the aggregated parameters. Assume that at each client $i$, a parameter denoted as $\theta_{i}$ is uploaded to a server for aggregating global parameter $\theta$. The procedure of secret aggregation is described in Algorithm 1. In the beginning, each client $C_{i}$ randomly generates parameters $\theta_{i}^{r}$ and shares it with other clients with encryption (Shown in Appendix A). Each client has a set of parameters $\{\theta_{1}^{r},\theta_{2}^{r}.\cdots,\theta_{n}^{r}\}$. For each client, $C_{i}$ uploads perturbed parameter $(\theta_{i}+\theta_{i}^{r})$ to server. The server collects all perturbed parameters from clients, computes aggregated perturbed parameter $\theta^{r}$, and sends it back to all clients. Finally, Clients can compute aggregated parameter $\theta$ and prevent from exposing it to the server.
0:  $n$ clients $C_{1},C_{2},\ldots,C_{n}$, client $C_{i}$ has parameter $\theta_{i}$
  Each client $C_{i}$ generates random parameter $\theta_{i}^{r}$
1  Transmit $\theta_{i}^{r}$ to all other clients with encryption, each
client has a set of parameters $\{\theta_{1}^{r},\theta_{2}^{r}.\cdots,\theta_{n}^{r}\}$
  Client $C_{i}$:
  Upload perturbed parameter $(\theta_{i}+\theta_{i}^{r})$ to server
  Receive parameters $\sum_{j=1}^{n}(\theta_{j}+\theta_{j}^{r})$ from server
  Compute parameters $\theta=[\sum_{j=1}^{n}(\theta_{j}+\theta_{j}^{r})-\sum_{i}^{n}\theta_{i}^{r}]/n$
  Server:
  for $i=1,\cdots,n$ do
     Receive parameters $(\theta_{i}+\theta_{i}^{r})$ from client $C_{i}$
  end for
  Compute parameters $\theta^{r}=\sum_{j=1}^{n}(\theta_{j}+\theta_{j}^{r})$
  Send parameters $\theta^{r}$ to all clients
Algorithm 1 Secret Aggregation
4.1.2. Model Training
Similar to (Chen et al., 2021), the server constructs a set of mapping matrices $\{\mathbf{M}^{i}\ \in\{0,1\}^{n\times n_{i}}\}_{i=1}^{N}$ and existence vectors $\{\mathbf{v}^{i}\in\{0,1\}^{n\times 1}\}_{i=1}^{N}$ to denote the entities in each client, where $n$ is the number of all unique entities in KG set and $n_{i}$ is the number of entities from client $C_{i}$. $\mathbf{M}^{i}_{m,n}=1$ if the $m$-th entity in entity table corresponds to the $n$-th entity from client $C_{i}$. $\mathbf{v}^{i}_{m}=1$ indicates that the $m$-th entity in entity table exists in client $C_{i}$.
FedCQA performs secret aggregation for entity embeddings. First, the client $C_{i}$ will randomly initialize the local entity embeddings $\mathbf{E}_{0}^{i}\in\mathbb{R}^{n_{i}\times d}$ and perturbation embeddings $\mathbf{E}_{r}^{i}\in\mathbb{R}^{n_{i}\times d}$. Every client will share the permutation embeddings with all clients. At round $t$, the server will select part of the clients $\mathbf{C}$ participating in the training. After local training of CQA on respective local knowledge graphs, client $C_{i}$ sends perturbed local entity embeddings $\mathbf{E}_{t}^{i}+\mathbf{E}_{r}^{i}$ to the server. The server will aggregate the entity embeddings using perturbed local embeddings:
(2) $\mathbf{E}_{t+1}^{r}\leftarrow\left(\mathbbm{1}\oslash\sum_{i\in\mathbf{C}}\mathbf{v}^{i}\right)\otimes\sum_{i\in\mathbf{C}}\mathbf{M}^{i}(\mathbf{E}_{t+1}^{i}+\mathbf{E}_{r}^{i}),$
[…]
(3) $\mathbf{E}_{t+1}^{r,i}\leftarrow{\mathbf{M}^{i}}^{\top}\mathbf{E}_{t+1}^{r},$
and the client $C_{i}$ can compute and update the local entity embeddings as:
(4) $\mathbf{E}_{t+1}^{i}\leftarrow\mathbf{E}_{t+1}^{r,i}-{\mathbf{M}^{i}}^{\top}\left(\mathbbm{1}\oslash\sum_{j\in\mathbf{C}}\mathbf{v}^{j}\right)\otimes\sum_{j\in\mathbf{C}}\mathbf{M}^{j}\mathbf{E}^{j}_{r}.$
After secret aggregation, the entity embeddings of all clients are shared without exposing the sensitive information to the server. Besides the training of entity embeddings, the operator networks are trained and aggregated using FedAvg (McMahan et al., 2017), and the detailed descriptions of the FedCQA are shown in Appendix B
4.2. Query Retrieval
After training, the server in FedCQA manages the process of retrieving answers to complex queries.
4.2.1. Query Encoding
Query encoding methods commonly represent queries to embeddings and retrieve answers according to scoring functions where similarity functions are widely used. FedCQA encodes queries and treats two types of queries differently. For in-graph queries, FedCQA directly encodes queries using corresponding local complex query answering models. For cross-graph queries, as the query involves entities from multiple knowledge graphs, we decompose the queries to atomic expressions and encode these atomic queries using local CQA models, and use global operator function models to compute the representations of the original queries.
4.2.2. Answer Retrieval
Because the entity embeddings are not stored in the central server, we can only retrieve answers from all distributed local knowledge graphs. Given a query encoding $q$, we score all the candidate entities at each local knowledge graph, at client $C_{i}$:
(5) $\mathbf{S}^{i}_{q}\leftarrow f^{i}_{s}(\mathcal{V}_{i},q)\in\mathbb{R}^{n_{i}\times 1},$
where $f^{i}_{s}$ is a score function in the client $C_{i}$. Then the score will be uploaded to the server to aggregate a score table for all unique entities in the knowledge graph sets:
(6) $\mathbf{S}\leftarrow\left(\mathbbm{1}\oslash\sum_{i=1}^{N}\mathbf{v}^{i}\right)\otimes\sum_{i=1}^{N}\mathbf{M}^{i}\mathbf{S}^{i}.$
The final answers to the queries are retrieved globally from the knowledge graph set according to the score table.
5. Experiments
In this section, we create a benchmark for distributed knowledge graph complex query answering problems and evaluate our proposed FedCQA’s performance on the benchmark.
5.1. Datasets and Experiment Setting
We introduce the detailed information of our used datasets and the setting of our experiments.
Figure 3. The query structures used for evaluation in the experiments. Naming for each query structure is provided under each subfigure, for brevity, the p, i, and u represent the projection, intersection, and union operations respectively.
5.1.1. Datasets
In our experiment, following previous work, we use the three commonly used knowledge graph datasets: FB15k (Bollacker et al., 2008; Bordes et al., 2013), FB15k-237 (Toutanova and Chen, 2015), and NELL995 (Carlson et al., 2010) to construct the distributed query answering benchmark. In each dataset, there are vertices describing entities and edges describing relations. To evaluate the distributed complex query, we conduct experiments assuming having 3 and 5 clients in each federated complex query answering system, respectively. We randomly select relations into clients and split triples into clients according to selected relations as a local knowledge graph. The triplets in each local knowledge graph are separated into training, validation, and testing with a ratio of 8:1:1 respectively. Following previous works (Ren et al., 2020), we construct training graph $\mathcal{G}_{train}$, validation graph $\mathcal{G}_{val}$, and test graph $\mathcal{G}_{test}$ in each client by training edges, training+validation edges, and training+validation+testing edges, respectively. The detailed statistics are listed in the table 1. #Clients denotes the number of clients, #Nodes, #Relations, #Edges denote the average number of nodes, relations, and edges in each client respectively.
Table 1. The statistics of three datasets used for experiments.
Graphs #Clients #Nodes #Relations #Edges
FB15k-237 3 13,651 79 103,359
5 12,639 47.4 62,015
FB15k 3 14,690 448.3 197,404
5 14,279 269 118,442
NELL995 3 40,204 66.7 47,601
5 28,879 40 28,560
5.1.2. Query Sampling
Following previous work (Hamilton et al., 2018; Bai et al., 2023b), we evaluate the complex query answering performance on the following eight general query types with abbreviations of $1p,2p,2i,ip,3i,pi,2u$, and $up$. As shown in figure 3, each subgraph denotes one query type, where each edge represents either a projection or a logical operator, and each node represents either a set of entities, the anchor entities and relations are to be specified to instantiate logical queries. We use the sampling method commonly used in previous works (Bai et al., 2023c; Ren et al., 2020) to randomly sample complex queries from knowledge graphs. We randomly sample two sets of queries from the knowledge graph sets: in-graph queries and cross-graph queries to evaluate local and global answer retrieval performance. For local model evaluation, we first obtain training, validation, and testing queries from the formerly constructed local knowledge graphs respectively. Then for the training queries, we conduct a graph search to find corresponding training answers on the local training graph. For the validation queries, we search for the answers on both the training graph and the validation graph and only use those queries that have different numbers of answers on the two graphs. For the testing queries, we use those queries that have different answers on the testing graph from answers on the training graph and validation graph. For global model evaluation, we construct global training, validation, and testing knowledge graphs using all local knowledge graphs, and sample testing queries with atomic expressions from different local knowledge graphs, finally we search for answers on three global knowledge graphs and only use those queries that have different answers on the testing graph from other two graphs. We collect statistics of complex queries in three datasets and the statistics are shown in Table 2. The number of in-graph queries is the average number of the client’s local queries.
Table 2. The statistics of queries sampled from three datasets used for experiments.
Graphs #C In-graph Cross-graph
Train. Valid. Test. Test.
FB15k-237 3 317,226 11,528 11,539 32,573
5 180,552 6,619 6,673 31,469
FB15k 3 592,573 19,206 19,267 53,660
5 344,418 11,409 11,437 53,154
NELL995 3 208,070 8,810 8,750 24,954
5 117,231 5,177 5,118 24,237
Table 3. The retrieval performance of distributed knowledge graph complex query answering models when there are 3 clients. The average results of HR@3 and MRR of all clients are reported. The best results are underlined. The best results of distributed models are in bold.
GQE Q2P tree-LSTM
In-graph Cross-graph In-graph Cross-graph In-graph Cross-graph
Graph Setting HR@3 MRR HR@3 MRR HR@3 MRR HR@3 MRR HR@3 MRR HR@3 MRR
Local 12.64 12.03 - - 14.55 13.63 - - 13.32 12.73 - -
[…]
Local 22.05 18.21 - - 24.32 22.64 - - 22.87 20.51 - -
[…]
FedR 20.29 18.61 - - 25.32 22.71 - - 23.64 20.97 - -
[…]
Local 11.85 11.03 - - 15.86 13.02 - - 13.85 13.85 12.94 -
[…]
FedR 12.01 11.23 - - 16.04 13.26 - - 12.48 11.67 - -
[…]
5.1.3. Baselines
We can use various existing query encoding methods as our local base model, to evaluate the effectiveness and generalization ability of our proposed FedCQA, we select three commonly used complex encoding methods GQE (Hamilton et al., 2018), Q2P (Bai et al., 2022), tree-LSTM (Bai et al., 2023c) as our base model. GQE is a graph query encoding model that encodes a complex query into a vector in embedding space; Q2P represents complex queries using multiple vectors; tree-LSTM recursively represents complex queries and treats all operations, entities, and relations as tokens.
To the best of our knowledge, there are no existing federated complex query answering methods but several federated knowledge graph embedding methods, therefore, we choose to compare our methods with FedE (Chen et al., 2021) and FedR (Zhang et al., 2022), two commonly used federated knowledge graph embedding methods as baselines. FedE aggregates both entity embeddings and relation embeddings in a server, while FedR only aggregates relation embeddings for privacy concerns and communication efficiency. We utilize these two methods with slight modifications to train a global complex query answering model: FedE aggregates all query encoding parameters and FedR aggregates relation embeddings and query encoding networks. Besides, we also compare our FedCQA with local and central settings. In the local setting, there is no collaboration between clients, while in the central setting, all distributed graphs in the knowledge graph set are aggregated for a global graph for training, we sample complex queries from global training, and validation knowledge graphs for training and validation.
If there is no further statement, we use the following implementation settings in the experiments. We evaluate FedCQA on three base models: GQE, Q2P, and tree-LSTM in FedE, FedR, and our proposed FedCQA for comparison. We tune hyper-parameters on the validation local queries for the base query encoding methods and set the dimension of entities and relations as 400 for all models for fair comparison and use AdamW (Loshchilov and Hutter, 2018) as optimizer.
Table 4. The retrieval performance of distributed knowledge graph complex query answering models when there are 5 clients. The average results of HR@3 and MRR of all clients are reported. The best results are in bold.
FB15k-237 FB15k NELL
In-graph Cross-graph In-graph Cross-graph In-graph Cross-graph
Graph Setting HR@3 MRR HR@3 MRR HR@3 MRR HR@3 MRR HR@3 MRR HR@3 MRR
Local 11.44 10.65 - - 14.65 13.8 - - 11.23 10.37 - -
GQE FedCQA 12.42 11.60 11.20 10.79 16.13 15.78 15.28 14.91 12.48 11.91 11.49 11.02
Local 19.83 17.51 - - 36.10 35.04 - - 20.03 18.62 - -
Q2P FedCQA 21.40 20.83 20.71 19.94 40.81 37.96 38.56 35.73 24.59 23.75 23.85 22.90
Local 10.48 10.09 - - 15.26 14.37 - - 14.52 13.89 - -
tree-LSTM FedCQA 13.79 13.27 12.74 12.18 15.44 15.81 15.28 14.24 15.68 14.28 14.57 12.89
(a) Cross-graph Queries
(b) Cross-graph Queries
(c) In-graph Queries
(d) In-graph Queries
Figure 4. The evaluation results of FedCQA-GQE facing different types of cross-graph queries on FB15k-237 (subfigures (a), (b)). The evaluation results of FedCQA-GQE model facing in-graph queries on FB15k-237 (subfigures (c), (d)).
5.1.4. Evaluation Metrics
Following the previous work (Bai et al., 2023b), we evaluate the generalization capability of models by calculating the rankings of answers that cannot be directly retrieved from an observed knowledge graph. Given a testing query $q$, the training, validation, and public testing answers are denoted as $M_{train}$, $M_{val}$, and $M_{test}$, respectively. We evaluate the quality of retrieved answers using Hit ratio (HR) and Mean reciprocal rank (MRR). HR@K metric evaluates the accuracy of retrieval by measuring the percentage of correct hits among the top K retrieved items. The MRR metric evaluates the performance of a ranking model by computing the average reciprocal rank of the first relevant item in a ranked list of results. The metric can be defined as:
(7) $\text{Metric}(q)=\frac{1}{|M_{test}/M_{val}|}\sum_{v\in M_{test}/M_{val}}m(rank(v)),\$
$m(r)=\textbf{1}[r\leq K]$ if the metric is HR@K and $m(r)=\frac{1}{r}$ if the metric is MRR. Higher values denote better reasoning performance. We train local models at each client by using the in-graph training queries and tune hyper-parameters using the validation queries. The evaluation is then finally conducted on the testing queries, including the evaluation of in-graph queries on local query encoding models and cross-graph queries on the global federated knowledge graph query answering system, respectively.
5.2. Performance Evaluation
We evaluate FedCQA’s complex query answering performance on three datasets and compare it to other baselines. We apply FedCQA on three base CQA models and evaluate the average retrieval performance on various queries. The results are summarized in Table 3 and Table 4. Table 3 reports the retrieval performance of various distributed knowledge graph complex query answering models when there are 3 clients. For each model, we evaluate performance facing in-graph queries and cross-graph queries, respectively. For in-graph queries, the average scores of all clients are reported. We report results in HR@3 and MRR which higher scores indicate better performance. The best results are underlined. The best results of distributed models are in bold. As shown in Table 3, our proposed methods can effectively retrieve complex query answers from distributed knowledge graphs. In comparison to local settings, we can see that FedCQA can utilize all participated knowledge graphs and performs better facing in-graph queries. For example, GQE model with FedCQA can retrieve 14.21 HR@3 on average while can only reach 11.85 without collaboration. Besides, in comparison to other federated knowledge graph embedding methods, our proposed FedCQA can reach comparable performance in both in-graph queries and cross-graph queries without exposing sensitive entity embeddings to the server. For example, FedCQA achieves the best performance in cross-graph queries in more than half of datasets and base CQA models.
In Table 4, we present the performance of FedCQA when there are 5 clients. We compare the performance with local training without collaboration to demonstrate the influence of client numbers. As shown in the table, FedCQA performs well compared to complex query answering models without collaboration, there are performance improvements in all datasets after applying FedCQA to various base CQA models, demonstrating that FedCQA can utilize the intrinsic information in the distributed knowledge sets. The collaboration in distributed knowledge graphs allows FedCQA to reason over various logical paths to improve performance.
5.3. Query Types
FedCQA can globally reason over distributed knowledge graphs and retrieve answers to cross-graph queries. To evaluate FedCQA’s performance on various types of complex queries, we conduct experiments to evaluate the retrieval performance of FedCQA and compare it to central learning on FB15k-237 when there are 3 clients. Figure 4(a) and Figure 4(b) show the FedCQA-GQE’s performance facing cross-graph queries. As we can see from the figure, FedCQA performs well on most various types of queries compared to the central model. For example, on query types ’2i’ and ’pi’, FedCQA can reach more than 90% MRR compared to the central model.
5.4. Local Influence
Because the in-graph queries can be processed by a single local knowledge graph CQA model, in this part, we evaluate the performance of FedCQA on these queries to assess the influence of FedCQA on local queries. We conduct experiments on FB15k-237 and the number of clients is 5. We evaluate the performance of FedCQA based on GQE and compare the model with no collaboration. The results are summarized in the Figure 4(c) and Figure 4(d). As shown in the figure, although each client has different performance due to the triplets correlation being different in each sub-dataset, FedCQA can improve all clients’ performance compared to the model without collaboration.
6. Conclusion
In this work, we present a federated complex query answering framework, FedCQA to reason over distributed knowledge sets with privacy preserved. We define the distributed knowledge graph complex query answering problem. To solve the problem, we propose secret aggregation for federated learning where the aggregated parameters can be kept secret to the server. Besides, we design a distributed query retrieval process for answering queries from distributed knowledge graph sets to protect clients’ privacy. To evaluate FedCQA model performance, we construct a benchmark based on three commonly used knowledge graph complex query answering datasets: FB15k-237, FB15k, and NELL. Extensive experiments on the benchmark demonstrate the effectiveness of our proposed FedCQA. FedCQA can retrieve answers given a query while keeping the sensitive information secret at local knowledge graphs. In the future, we aim to propose new methods for better answering complex queries by exploiting intrinsic information in the distributed knowledge graphs.
ACKNOWLEDGMENTS
The authors of this paper were supported by the NSFC Fund (U20B2053) from the NSFC of China, the RIF (R6020-19 and R6021-20) and the GRF (16211520 and 16205322) from RGC of Hong Kong. We also thank the support from UGC Research Matching Grants (RMGS20EG01-D, RMGS20CR11, RMGS20CR12, RMGS20EG19, RMGS20EG21, RMGS23CR05, RMGS23EG08).
Appendix A Parameter Sharing
In this section, we provide an example of sharing secrets between clients. The parameters can be shared under the protection using various encryption methods, for example, the commonly used is Diffie–Hellman key exchange (Diffie and Hellman, 2022) shown as follows, we consider the sharing process between two clients:
  • Client A and client B publicly agree to use a modulus $p$ and base $g$, $p$ is a prime.
  • Client A chooses a secret integer $a$, then sends client B $m_{A}=g^{a}\mod p$.
  • Client B chooses a secret integer $b$, then sends Alice $n_{B}=g^{b}\mod p$.
  • Client A computes $s=m_{B}^{a}\mod p$.
  • Client B computes $s=m_{A}^{b}\mod p$.
After D-H key exchange, Client A and B share a secret $s=g^{ab}\mod p$. The secret $s$ can be used as encryption to share sensitive information between clients.
Appendix B Algorithm
In this section, we present the pseudo code for FedCQA.
10:  The number of clients $N$; The faction of clients
selected in each round $F$;
  Client $C_{i}$:
  Client $C_{i}$ initialize entity embeddings $\mathbf{E}^{i}$ and perturbation embeddings $\mathbf{E}_{r}^{i}$.
  Share $\mathbf{E}_{r}^{i}$ with other clients with encryption
  Receive and decryption to get $\{\mathbf{E}_{r}^{1},\cdots,\mathbf{E}_{r}^{N}\}$
  Upload $\mathbf{E}^{i}$ to server for secret aggregation, receive $\mathbf{E}_{0}^{i}$
  Server:
  Server constructs permutation matrices $\{M_{i}\}_{i=1}^{N}$, and existence vectors $\{v_{i}\}_{i=1}^{N}$ and initialize operator networks $\Theta_{0}$, distribute to all clients.
  for $t=0,1,2,\cdots$ do
     Server distributes operator networks to each client.
     $\mathcal{C}_{t}$leftarrow Randomly select client set with $N\times F$ clients.
     for $C_{i}\in\mathcal{C}_{t}$ in parallel do
        $(\mathbf{E}^{i}_{r}+\mathbf{E}^{i}_{t+1}),\Theta^{i}_{t+1}\leftarrow\textbf{ClientUpdate}(C_{i},{\mathbf{M}^{i}}^{\top}\mathbf{E}^{r}_{t},\Theta_{t})$
     end for
     $\mathbf{E}_{t+1}^{r}\leftarrow\left(\mathbbm{1}\oslash\sum_{i\in\mathbf{C}_{t}}\mathbf{v}^{i}\right)\otimes\sum_{i\in\mathbf{C}_{t}}\mathbf{M}^{i}(\mathbf{E}_{t+1}^{i}+\mathbf{E}_{r}^{i})$
     $\Theta_{t+1}\leftarrow 1/|\mathcal{C}_{t}|\sum_{i\in\mathcal{C}_{t}}\Theta_{t}^{i}$
  end for
[…]
  $\mathbf{E}\leftarrow\mathbf{E}^{r}-{\mathbf{M}^{i}}^{\top}\left(\mathbbm{1}\oslash\sum_{j\in\mathbf{C}}\mathbf{v}^{j}\right)\otimes\sum_{j\in\mathbf{C}}\mathbf{M}^{j}\mathbf{E}^{j}_{r}$
  for $e=1,\cdots,E$ do
     $\mathbf{E},\Theta\leftarrow\textbf{LocalUpdate}(\mathbf{E},\Theta)$
  end for
  return  $\mathbf{E}+\mathbf{E}^{i}_{r},\Theta$
Algorithm 2 FedCQA Framework